Grade 11Vectors

What is the dot product of two vectors of magnitude 3 and 5 if the angle between them is 60Degree

Rohit Kumar
8 Years agoGrade 11

1 Answer

Om
8 Years ago
Dot product of two vectors is satisfied by function of costheta. So3 • 5 = 3×5cos60 =15*1\2 =7•5And for cross product use sintheta